OpenStudy (anonymous):

If 2.0 liters of oxygen gas react with excess carbon monoxide at STP, how many liters of carbon dioxide can be produced under the same conditions? 2 CO (g) + O2 (g) 2 CO2 (g)

OpenStudy (anonymous):

here u see coefficeint of O2 is 1 and coefficeint of CO2 is 2 so when u get the mole of O2 u just multiply it by 2 u will get ans as n=v/22.4 as u know R=0.082 and temperature at STP is 298K ok

OpenStudy (anonymous):

so PV=nRT and at STP always remember RT=22.4 if u are given temp. 27C then RT=24.6 P=i atm u ll get n=V/RT

OpenStudy (anonymous):

u done?

OpenStudy (anonymous):

Yes! thank you

OpenStudy (anonymous):

yw :)